### v2

# Concept Based Continuous Prompts for Interpretable Text Classification

Abstract
Continuous prompts have become widely adopted for augmenting performance across a wide range of natural language tasks. However, the underlying mechanism of this enhancement remains obscure. Previous studies rely on individual words for interpreting continuous prompts, which lacks comprehensive semantic understanding. Drawing inspiration from Concept Bottleneck Models, we propose a framework for interpreting continuous prompts by decomposing them into human-readable concepts. Specifically, to ensure the feasibility of the decomposition, we demonstrate that a corresponding concept embedding matrix and a coefficient matrix can always be found to replace the prompt embedding matrix. Then, we employ GPT-4o to generate a concept pool and choose potential candidate concepts that are discriminative and representative using a novel submodular optimization algorithm. Experiments demonstrate that our framework can achieve similar results as the original P-tuning and word-based approaches using only a few concepts while providing more plausible results.
Concept Based Continuous Prompts for Interpretable Text Classification
Qian Chen1, Dongyang Li1, Xiaofeng He1,2 thanks: Correspondingauthor 1 School of Computer Science and Technology, East China Normal University, Shanghai, China 2 NPPA Key Laboratory of Publishing Integration Development, ECNUP, Shanghai, China {qianchen901005,fromdongyang}@gmail.com, hexf@cs.ecnu.edu.cn
1 Introduction
Continuous prompts are widely used in NLP with stable performance and easy-hand training compared to discrete prompts. However, the introduction of dense prompt tokens compromises the interpretability.
Recent trials exploring the underlying mechanism found that models do not rely on context augmented by complementary prompt tokens, which is puzzling Kavumba et al. (2022). Hambardzumyan et al. (2021) suggest that adding non-perceptible perturbation to prompt embedding could bring a different outcome. Khashabi et al. (2022) propose Prompt Waywardness Hypothesis, which means it is impossible to match a continuous prompt with one exact discrete text. Ju et al. (2023) introduce numerous words to mitigate the gap between discrete tokens and dense learned embedding. However, generated word explanations may lack comprehensible and integral semantics. For example, in Figure 1, although given top-5 selected words, they are just listed without organic combination, being blended with stop words noise.
Figure 1: An explanation example from Ju et al. (2023). From left to right three sub-figures represent the SST-2, IMDB, and AGNews dataset results. The color bar represents values.
Moreover, due to the infeasibility of discrete words regression to continuous prompt, the inherent gap would driven the abuse of words, as shown in Figure 2, simply increasing vocabulary capacity fails to rise up the performance. And word-based methods cannot apply to BPE encoding Sennrich et al. (2016) based models directly like GPTs Radford et al. (2018, 2019).
Inspired by Concept Bottleneck Models (CBMs) Koh et al. (2020), we propose Concept Decomposition (CD) framework, which decomposes the prompt embeddings into concepts. In detail, we start by reclaiming the formulation of the vanilla transformer block, in which we find that the outputs essentially correspond to linear combinations of column vectors from the outer layer of the Feed-Forward Network (FFN). Based on findings in Geva et al. (2021), we argue that continuous prompts could be decomposed into several comprehensible fragmented semantic patterns, which can be aligned with integral human-readable concepts. In order to complete the pattern-concept alignment, we first utilize the powerful GPT-4o OpenAI et al. (2024) to construct an ample space of possible concepts owing to Large Language Models (LLMs) containing significant world knowledge that can be elicited by instructions instead of labor-intensively annotation Geva et al. (2021).
After construction, as we require highly discriminative concepts that cover various aspects to obtain a better explanation, we formulate above requirements into a submodular optimization problem that allows us to choose good concepts efficiently Nemhauser et al. (1978a).
Figure 2: Left: The bad case from Ju et al. (2023), which the continuous prompt $\boldsymbol{p}$ fails to locate in the span space $\boldsymbol{V}$. Given the existence of $\boldsymbol{p}_{1}$, the fitting gap cannot be eliminated. Right: The accuracy results from Ju et al. (2023). x represents a vocabulary size. As the capacity of vocabulary increases, the performance does not significantly change.
Extensive experiments demonstrate the effectiveness of our framework. In the evaluation of two typical model architectures (encoder-only and decoder-only) on three diverse datasets, CD achieves comparable performance to the original continuous prompt method both in few-shot and full-shot scenarios, providing more meaningful and comprehensible concepts with superior to the discrete explanation method.
Our contributions are summarized as follows:
  • We prove that one can always find a concept embedding matrix and a coefficient matrix corresponding to the continuous prompt matrix.
  • We propose a framework (CD) to maintain representative concepts for interpreting continuous prompts.
  • We evaluate the proposed framework on three classification datasets with two typical language models: BERT Devlin et al. (2019) and GPT-2 Radford et al. (2019), and the results demonstrate the effectiveness.
2 Related Work
2.1 Concept Bottleneck Models
CBMs have garnered significant attention, with specific concepts serving as fundamental units of explanation, also known as rationales. These models solve for rationales and combine them to produce outputs. For instance, Yeh et al. (2020) propose a concept discovery method that aims to infer a complete set of concepts. Recently, CBMs have been found to have a backdoor path due to inherent structures Kumar and Talukdar (2020), which complicates their integration with causal inference. Zhang et al. (2023) propose two casual metrics to identify valid rationales when two or more snippets are highly inter-correlated.
2.2 Continuous Prompts and Explanations
Continuous prompts refer to continuous vectors incorporated into the input of a prompted language model. These prompts are typically tuned using gradient-based methods, which are directed by the labeled training examples of the tasks Qin and Eisner (2021); Sučik et al. (2023); Huang et al. (2023); Shah et al. (2023); Liu et al. (2023). Although these prompts generally enhance model performance, their continuous nature poses challenges for human understanding and interpretation Kavumba et al. (2022); Hambardzumyan et al. (2021); Khashabi et al. (2022); Ju et al. (2023). Hambardzumyan et al. (2021) test whether few-shot prompt-based models depend on shallow cues, the results reveal that simple shortcuts still exist. Khashabi et al. (2022) use adversarial prompts to inject knowledge into models, outperforming manually designed prompts. Ju et al. (2023) propose to use discrete words to explain the continuous prompts in a bag-of-words manner. The interpretations, as shown in Figure 1, made up of words sorted by the learned weights, exhibit fragmented semantics, which impairs plausibility.
3 Start From Attention
In this section, we give a brief formulation on how the continuous prompts can be decomposed into semantically related concepts. Given a prompt augmented sequence of token representations $\boldsymbol{T}=(\boldsymbol{t}_{1},\cdots,\boldsymbol{t}_{n})\in\boldsymbol{R}^{d\times n}$, and a model with $H$ heads. $d$ is the embedding dim and $d_{h}=d/H$. For simplicity, we suppose the continuous token is the $i$-th token. Each Multi-Head Attention (MHA) computes $\boldsymbol{z}_{i}^{h}$ as follows:
$\boldsymbol{z}_{i}^{h}=\sum_{j}\boldsymbol{A}_{i,j}^{h}\boldsymbol{W}_{V}^{h}\boldsymbol{t}_{j}$ (1)
with $\boldsymbol{A}_{i,j}^{h}$ referring to the attention weight where token $i$ attends token $j$, and $\boldsymbol{W}_{V}^{h}\in\mathbb{R}^{d_{h}\times d}$ is the weight matrix of $h$-th head.
Generally, the FFN block has two layers, and we use $\boldsymbol{r}_{m}$, $\boldsymbol{p}_{m}\in\mathbb{R}^{d}$ to denote the $m$-th column vectors of the inner weight matrix $\boldsymbol{W}_{1}\in\mathbb{R}^{d\times M}$ and outer weight matrix $\boldsymbol{W}_{2}\in\mathbb{R}^{d\times M}$. Suppose the activation function is $\phi$. We can obtain the final output $\boldsymbol{y}_{i}$ as follows:
$\begin{split}\boldsymbol{y}_{i}&=\sum_{h}\sum_{j}\sum_{m}\boldsymbol{A}_{i,j}^{h}\phi(\boldsymbol{r}_{m}^{T}\boldsymbol{W}_{O}^{h}\boldsymbol{W}_{V}^{h}\boldsymbol{t}_{j})\boldsymbol{p}_{m}+\\ &\sum_{h}\sum_{j}\sum_{m}\phi(\boldsymbol{r}_{m}^{T}\boldsymbol{t}_{j})\boldsymbol{p}_{m}\end{split}$ (2)
with $m$ referring the $m$-th dimension of inner hidden states. $\boldsymbol{W}_{O}^{h}\in\mathbb{R}^{d\times d_{h}}$ is the partitioned weight matrix. The second term is deduced by the residual connection trick. As the the Layer Normalization can be reformulated into a linear transformation operation Ferrando et al. (2022), we only take the residual into consideration.
Notice that all the terms preceding $\boldsymbol{p}_{m}$ are scalars, so the final output can be viewed as a linear combination of $\boldsymbol{p}_{m}$. Based on findings of Geva et al. (2021), each $\boldsymbol{p}_{m}$ corresponds to a shallow or semantic pattern, which can be described with concepts. Therefore, we decompose the continuous prompts into human-readable concepts.
4 Problem Formulation
Given a training set of text $\mathcal{X}=\{(x,y)_{l}\mid l\in|\mathcal{X}|\}$, where $x$ is the text and $y\in\mathcal{Y}$ represents a label from a set of $N$ classes. Normally we use $\mathcal{X}$ to tune the continuous prompts $\boldsymbol{P}\in\mathbb{R}^{d\times N_{q}}$ with $N_{q}$ being the number of continuous prompt tokens. Let $\mathcal{S}$ be the superset of candidate textual concepts generated from GPT-4o and $\mathcal{C}\subseteq\mathcal{S}$ be the desired set with $N_{c}$ concepts.
We then initialize concept embeddings $\boldsymbol{C}\in\mathbb{R}^{d\times N_{c}}$ and coefficient $\boldsymbol{Q}\in\mathbb{R}^{N_{c}\times N_{q}}$. Our goal is to decompose the continuous prompts into $\boldsymbol{CQ}$ and ensure the explanation fidelity Yeh et al. (2019).
Now we give the following lemma to ensure the decomposition is feasible [^1] The proof is available in Appendix A..
Lemma For any $\boldsymbol{P}\in\mathbb{R}^{d\times N_{q}}$, $\epsilon>0$, there exist $\boldsymbol{C}\in\mathbb{R}^{d\times N_{c}}$, $\boldsymbol{Q}\in\mathbb{R}^{N_{c}\times N_{q}}$ that satisfies $||\boldsymbol{CQ}-\boldsymbol{P}||_{F}^{2}\leq\epsilon$.
With this lemma, we can assure that for each continuous prompts, a corresponding concept can be identified.
5 Methodology
Figure 3: Overview of our framework CD. Our framework prompts GPT-4o to generate candidate concepts for each class (Section 5.1). We then use submodular optimization to select ones that maximize diversity and coverage (Section 7). Next, we initialize the concept embedding $\boldsymbol{C}$ by feeding the concepts into a text encoder (BERT-Large or GPT2-Medium) and initialize the coefficient embedding $\boldsymbol{Q}$. Finally, we tune the embeddings using stochastic gradient descent (Section 5.3). $\tilde{\boldsymbol{X}}$ denotes the prompt-augmented input. $\tilde{\boldsymbol{P}}=\boldsymbol{CQ}$.
As illustrated in Figure 3, our framework prompts GPT-4o to generate a set of candidate concepts for each class (Section 5.1). We employ submodular optimization to greedily select a subset of concepts for each class such that we maximize diversity and coverage scores (Section 7). We then initialize the concept embedding $\boldsymbol{C}$ by feeding the concepts into a text encoder and the coefficient embedding $\boldsymbol{Q}$. Ultimately we design a loss to tune the two embeddings (Section 5.3).
5.1 Generate the Candidates
We first generate a candidate set of concepts by combining prompts from labeled text. To ensure text diversity, we generate 50 samples and design corresponding prompts for the experimental dataset, as shown in Table 1. Universal prompt templates need to adapt to specific information in the dataset. For instance, in the IMDB dataset Maas et al. (2011), template 1 will be modified to “describe what a positive review looks like”. Additionally, due to the uncontrollability of LLMs Bhargava et al. (2024), we append auxiliary instructions following prompts to avoid label leakage. However, even with these precautions, we still find some text instances containing labels. Therefore, we remove parts of the generated text through regular expression matching.
Universal Prompt Template
1. describe what a [CLASS NAME] looks like:
2. describe the aspects of a [CLASS NAME]:
3. describe [CLASS NAME] using sentences:
4. describe the patterns of a [CLASS NAME]:
5. describe the concepts of a [CLASS NAME]:
Table 1: The prompt templates used to generate raw sentences from GPT-4o.
5.2 Submodular Optimization
To ensure that each category contributes equally to each concept, we need to select $k$ relevant concepts for each category $y$. We use the corresponding prompts from Table 1 to generate the overall candidate set $\mathcal{S}$, and we split $\mathcal{S}$ for each category as $\mathcal{S}_{y}$.
The next step is to select the desired set $\mathcal{C}_{y}$ from $\mathcal{S}_{y}$, which satisfies two properties: diversity and coverage. Diversity means that concepts are more informative, while the second requires good coverage of the candidate set. This problem is a typical combinatorial optimization problem. Fortunately, we can solve it in polynomial time complexity using submodular optimization Iyer and Bilmes (2013). We can measure the two properties above with a function $\mathcal{F}:2^{|\mathcal{C}_{y}|}\rightarrow\mathbb{R}$. A submodular function satisfies the diminishing property [^2] A set function $\mathcal{F}$ satisfying diminishing property is that for every $\mathcal{A}\subseteq\mathcal{B}$ and $x\notin\mathcal{B}$, $\mathcal{F}(\mathcal{A}\cup\{x\})-\mathcal{F}(\mathcal{A})\geq\mathcal{F}(\mathcal{B}\cup\{x\})-\mathcal{F}(\mathcal{B}).$. If the diminishing function is monotonic [^3] A monotonic submodular function $\mathcal{F}$ is that for every $\mathcal{A}\subseteq\mathcal{B},\mathcal{F}(\mathcal{A})\leq\mathcal{F}(\mathcal{B}).$, then the problem can be solved using a greedy algorithm Nemhauser et al. (1978b). Inspired by Lin and Bilmes (2011), we propose the following monotonic submodular function to select $\mathcal{C}_{y}$ from $\mathcal{S}_{y}$:
$\mathcal{F}(\mathcal{C}_{y})=\lambda\mathcal{L}_{d}(\mathcal{C}_{y})+\mathcal{L}_{c}(\mathcal{C}_{y})$ (3)
with $\mathcal{L}_{d}(\mathcal{C}_{y})$ evaluates diversity and $\mathcal{L}_{c}(\mathcal{C}_{y})$ rewards coverage. The first term would give a higher score when containing informative concepts, and the second term is higher when concepts are representative. The hyperparameter $\lambda$ controls the weights of two terms.
Diversity Score We propose the following diversity score:
$\mathcal{L}_{d}(\mathcal{C}_{y})=\sum_{c\in\mathcal{C}_{y}}\psi\left(\sum_{y\in\mathcal{Y}}Div(y,c)\right)$ (4)
with $\psi$ is the concave function and $Div(y,c)$ measures the differences between $\mathcal{C}_{y}$ and $\mathcal{Y}$. Here, we use a simple affine function (i.e. $\psi(x)=x$). And we calculate $Div(y,c)$ as follows:
$Div(y,c)=\sum_{y\in\mathcal{Y}}Sim(y|c)\cdot\text{log}\left(Sim(y|c)\right)$ (5)
$Sim(y,c)=\frac{1}{|\mathcal{X}_{y}|}\sum_{x\in\mathcal{X}_{y}}\mathcal{E}(x)\cdot\mathcal{E}(c)^{T}$ (6)
with $\mathcal{E}(\cdot)$ represents the encoder.
Coverage Score We define the following coverage score:
$\mathcal{L}_{c}(\mathcal{C}_{y})=\sum_{c_{1}\in\mathcal{S}_{y}}\sum_{c_{2}\in\mathcal{C}_{y}}\phi(c_{1},c_{2})$ (7)
where $\phi$ represents the similarity measure between $c_{1}$ and $c_{2}$. This score is a minimax facility location metric Ziaei et al. (2013), aiming to minimize the maximum distance between each element in the subset and the candidate set. The distance measure employed is the cosine similarity between the features of the two concepts extracted by the text encoder. A high coverage score indicates a diverse concept encompassing a target class’s potential aspects.
5.3 Optimize Interpretations
In this section, we introduce the method for solving $\boldsymbol{C}$ and $\boldsymbol{Q}$. A classic post-hoc explanation system derives the corresponding explanation based on the input and output Kenny et al. (2021). We require our explanations to stay consistent with original continuous prompts while maintaining the performance. Based on prior studies Yeh et al. (2019); Ju et al. (2023), we construct two loss terms to meet these requirements. For mitigating deviation, we introduce the following constraint:
$\mathcal{L}_{f}=\mathcal{D}_{\text{KL}}(\mathcal{P}(y|\boldsymbol{X},\boldsymbol{P})\parallel\mathcal{P}(y|\boldsymbol{X},\boldsymbol{CQ}))$ (8)
where $\mathcal{P}(\cdot)$ represents the output probability. For performance retention, we introduce the following constraint:
$\mathcal{L}_{l}=-\text{log}\mathcal{P}(y|\boldsymbol{X},\boldsymbol{CQ}).$ (9)
We combine these two terms in the following manner:
$\mathcal{L}=\mu\mathcal{L}_{f}+\mathcal{L}_{l}$ (10)
with $\mu\in[0,1]$.
5.4 Local Level Interpretations
In this section, we deliberately introduce our strategy for generating explanations. Explaining a continuous prompt is non-trivial. The difficulty lies in the fact that a global explanation may not be plausible enough Ju et al. (2023). We are more concerned with why a continuous prompt model makes a particular decision given the input and output (i.e. post-hoc manner).
Therefore, we propose the following local level explanation strategy. For input $(x,y)_{l}$, we optimize $\mathcal{F}(\mathcal{C}_{y})$ (3) to obtain candidates. Then we select top-k concepts as our explainable concepts sorted by keys derived from $\boldsymbol{Q}$. The sorting key of $c_{y,p}$ is calculated as:
$\text{key}(c_{y,p})=\sum_{q=0}^{N_{q}-1}\boldsymbol{Q}_{p,q}$ (11)
where $c_{y,p}$ denotes the $p$-th element of $\mathcal{C}_{y}$.
6 Experiments
6.1 Experimental Setups
Metrics and Baselines We investigate the interpretability of CD by comparing it to other explanation methods using two metrics: accuracy and concept correlation. Following Ju et al. (2023), we use accuracy to assess how well the concept prompt aligns with the continuous prompts. We adopt Vocab-1500 explanation methods as our baseline, with 1500 being the maximum number of vocabularies according to Ju et al. (2023). Inspired by Jin et al. (2020), we introduce a new metric called Concept Correlation, denoted as $\rho$, to evaluate concept-level explanations. Concretely, we compute the Pearson correlation between the coefficients learned by a bag-of-concepts model and the attribution scores provided by various explanation methods. For baselines, we choose three attribution methods to assign scores for each prompt. These scores are comparable to coefficients for explanation faithfulness. Specifically we choose TokenSHAP Goldshmidt and Horovicz (2024), IG Sundararajan et al. (2017), and grad Shrikumar et al. (2017).
Datasets and Backbones Throughout our experiments, we primarily study BERT-Large (335M) and GPT-2-Medium (345M) on text classification tasks. In Section 6.2, we utilize the SST-2 Socher et al. (2013), IMDB Maas et al. (2011), and AGNews Zhang et al. (2015) datasets.
Implementation Details We implement both full and few-shot learning settings (i.e. 4-shot, 8-shot, 16-shot, 32-shot, and full-shot). We design two different types of prompts for each dataset and customize them to suit diverse datasets [^4] Appendix F provides results with more types of prompts.. Throughout our experiments, we implement four different prompt length settings, including 1, 2, 5, 10. Following Liu et al. (2023), we train continuous prompts and concept embeddings using the AdamW optimizer with a learning rate of $10^{-4}$. We employ early stopping with 100 iteration steps. Additional hyperparameters are provided in the Appendix. Results are reported using five random seeds.
BERT-Large GPT-2-Medium
$N_{q}=1$ $N_{q}=2$ $N_{q}=1$ $N_{q}=2$
Dataset   shot P-tuning CD Vocab P-tuning CD Vocab P-tuning CD P-tuning CD
  SST-2 4 0.581 0.592 0.556 0.625 0.606 0.620 0.560 0.557 0.533 0.545
8 0.691 0.615 0.618 0.579 0.674 0.642 0.636 0.596 0.605 0.603
16 0.840 0.797 0.818 0.744 0.744 0.752 0.831 0.848 0.821 0.809
32 0.847 0.850 0.859 0.831 0.812 0.782 0.839 0.852 0.832 0.856
Full 0.910 0.907 0.907 0.896 0.892 0.904 0.905 0.917 0.906 0.840
Avg 0.773 0.752 0.751 0.735 0.745 0.740 0.754 0.754 0.739 0.730
${\sigma^{2}}$${\pm 0.028}$${\pm 0.005}$${\pm 0.003}$$\pm 0.002$$\pm 0.003$$\pm 0.004$$\pm 0.0004$$\pm 0.0005$$\pm 0.0008$$\pm 0.003$
  IMDB 4 0.542 0.577 0.556 0.563 0.538 0.569 0.547 0.556 0.529 0.529
8 0.750 0.789 0.688 0.605 0.549 0.536 0.509 0.507 0.518 0.501
16 0.834 0.876 0.852 0.848 0.772 0.759 0.506 0.507 0.504 0.527
32 0.839 0.848 0.867 0.804 0.800 0.800 0.790 0.714 0.763 0.864
Full 0.918 0.917 0.919 0.919 0.916 0.917 0.924 0.926 0.909 0.919
Avg 0.776 0.801 0.776 0.747 0.715 0.716 0.655 0.642 0.644 0.668
$\sigma^{2}$$\pm 0.004$$\pm 0.0009$$\pm 0.002$$\pm 0.001$$\pm 0.003$$\pm 0.004$$\pm 0.0004$$\pm 0.001$$\pm 0.001$$\pm 0.0005$
  AGNews 4 0.687 0.704 0.712 0.683 0.698 0.663 0.762 0.755 0.759 0.762
8 0.742 0.692 0.689 0.712 0.711 0.741 0.790 0.790 0.797 0.796
16 0.767 0.766 0.758 0.759 0.755 0.747 0.784 0.793 0.781 0.804
32 0.823 0.814 0.828 0.828 0.837 0.805 0.836 0.837 0.831 0.831
Full 0.882 0.880 0.884 0.888 0.887 0.885 0.865 0.866 0.854 0.858
Avg 0.780 0.771 0.774 0.774 0.777 0.768 0.807 0.808 0.804 0.810
$\sigma^{2}$$\pm 0.0009$$\pm 0.0006$$\pm 0.001$$\pm 0.0006$$\pm 0.0007$$\pm 0.001$$\pm 0.00003$$\pm 0.00005$$\pm 0.0001$$\pm 0.0001$
Table 2: Accuracy results of CD with baselines on the SST-2, IMDB and AGNews dataset. $N_{q}$ denotes the number of continuous prompt tokens. Vocab represents Vocab-1500. $\sigma^{2}$ denotes the average variance across five shot settings.
     BERT-Large GPT-2-Medium
Dataset top-k CD TokenSHAP IG Grad CD TokenSHAP IG Grad
  SST-2 3 0.755 0.646 0.525 0.481 0.824 0.613 0.535 0.342
5 0.561 0.427 0.221 0.264 0.610 0.497 0.353 0.289
10 0.484 0.398 0.141 0.173 0.445 0.310 0.252 0.200
Avg 0.600 0.490 0.295 0.306 0.626 0.473 0.380 0.277
$\sigma^{2}$$\pm 0.046$$\pm 0.034$$\pm 0.062$$\pm 0.038$$\pm 0.062$$\pm 0.041$$\pm 0.017$$\pm 0.029$
  IMDB 3 0.553 0.528 0.669 0.443 0.824 0.557 0.466 0.382
5 0.654 0.535 0.429 0.294 0.610 0.476 0.302 0.235
10 0.378 0.203 0.184 0.217 0.445 0.310 0.220 0.187
Avg 0.528 0.422 0.427 0.318 0.626 0.447 0.329 0.268
$\sigma^{2}$$\pm 0.039$$\pm 0.025$$\pm 0.038$$\pm 0.028$$\pm 0.062$$\pm 0.026$$\pm 0.024$$\pm 0.033$
  AGNews 3 0.553 0.505 0.629 0.557 0.824 0.513 0.299 0.432
5 0.654 0.454 0.424 0.333 0.610 0.496 0.320 0.297
10 0.378 0.219 0.203 0.155 0.445 0.217 0.156 0.309
Avg 0.528 0.392 0.418 0.348 0.626 0.408 0.258 0.346
$\sigma^{2}$$\pm 0.059$$\pm 0.034$$\pm 0.047$$\pm 0.029$$\pm 0.059$$\pm 0.077$$\pm 0.161$$\pm 0.027$
Table 3: Concept correlation results of CD with baselines. $\sigma^{2}$ denotes the average variance across three top-k settings.
  Class Input Top-3 Concepts
world euters china has said no date has been set for working level talks on the north korean nuclear crisis … 1.Breaking news updates on significant global events. 2.In-depth analysis of international conflicts and resolutions. 3.Coverage of diplomatic relations between countries.
sports michael phelps won the gold medal in the 400 individual medley … 1.Live coverage and updates of ongoing sporting events. 2.In-depth analysis of match statistics and player performances. 3.Profiles of athletes and their journeys to success.
business new york reuters u s treasury debt prices slipped on monday … 1.In-depth analysis of stock market trends and movements. 2.Coverage of corporate earnings reports and financial performance. 3.Updates on global economic indicators such as GDP and inflation.
sci/tech a company … won a grant to develop a method of producing better peptides … 1.Breaking news on scientific discoveries and breakthroughs. 2.In-depth analysis of emerging technologies and their implications. 3.Profiles of leading scientists, engineers, and innovators.
Table 4: Explanation examples on AGNews dataset with BERT-Large. Top-3 concepts are ranked based on the coefficients in $\boldsymbol{Q}$.
6.2 General Results
We first evaluate CD using accuracy metric across three datasets, as shown in Table 2. In summary, we can draw following two conclusions: i) CD could reach a similar level to P-tuning (e.g. BERT-Large/$N_{q}=1$, P-tuning: 0.773, CD: 0.752), which demonstrates our effectiveness in fitting continuous prompts with two typical language model architectures, further reflecting fidelity. The subtle differences demonstrate the feasibility of our framework. ii) CD, with concepts whose average lengths reach up to 300, can perform similarly to the discrete explanation method with 1500 vocabulary capacity (e.g. BERT-Large/$N_{q}=2$, CD:0.715, Discrete:0.716). Moreover, in some settings outperforms Discrete (e.g. BERT-Large/IMDB/$N_{q}=1$, CD: 0.801, Discrete: 0.776).
We then conduct correlation experiments on these three datasets with the full shot setting, and the results are shown in Table 3. We find that: i) In 7 out of 9 experimental settings, we achieved the best results, indicating that our method is faithful to continuous prompts. ii) As the number of concepts increases, the correlation decreases rapidly, which suggests that our explanations still suffer from noise issues. iii) We test different continuous prompt tokens and find that no significant difference, which demonstrates the generalizability of out method. We conduct more depth analysis about trade offs between interpretability and performance of CD in Section C.
Figure 4: Comparison results before and after attack on SST-2, IMDB and AGNews datasets on BERT-Large.
BERT-Large GPT-2-Medium
$N_{q}=5$ $N_{q}=10$ $N_{q}=5$ $N_{q}=10$
P-tuning CD Discrete P-tuning CD Discrete P-tuning CD P-tuning CD
  4-shot 0.519 0.588 0.533 0.543 0.514 0.592 0.551 0.542 0.542 0.553
8-shot 0.514 0.525 0.563 0.583 0.615 0.524 0.620 0.597 0.628 0.583
16-shot 0.646 0.773 0.682 0.746 0.797 0.790 0.792 0.804 0.769 0.772
32-shot 0.776 0.833 0.795 0.793 0.850 0.804 0.822 0.823 0.814 0.841
Full 0.868 0.897 0.874 0.909 0.883 0.869 0.894 0.906 0.894 0.908
Avg 0.664 0.723 0.689 0.714 0.731 0.715 0.735 0.734 0.729 0.731
${\sigma^{2}}$${\pm 0.017}$${\pm 0.009}$${\pm 0.005}$$\pm 0.001$$\pm 0.007$$\pm 0.008$$\pm 0.0009$$\pm 0.0007$$\pm 0.0006$$\pm 0.004$
Table 5: Accuracy results of CD with baselines on the SST-2 dataset. $\sigma^{2}$ denotes the average variance across five shot settings.
6.3 Interpretation Visualization
  Top-3 Concepts
1.Breaking news on mergers and acquisitions. 2.In-depth analysis of stock market trends and movements. 3.Profiles of successful entrepreneurs and leaders.
Table 6: One business news from AGNews dataset. The input is london reuters oil prices surged to a new high of 47 a barrel on wednesday after a new threat by rebel militia against iraqi oil facilities and as the united states said inflation had stayed in check despite rising energy costs.
According to results from Table 3, we choose top-3 concepts as sample explanations because top-5 or top-10 concepts contain noises which decreases the correlations. The examples are provided in the Table 4. As seen, our generated explanations align well with the input, resulting in credible explanations.
Nonetheless, there still exists some confusing explanations which the correspondence is not obvious. Take Table 6 as an example, the input is about oil price fluctuating while the first concept focuses on company merge. One possible reason may be that continuous prompts rely on short cut links Hambardzumyan et al. (2021); Ju et al. (2023). We conjecture that one reason of that maybe short cut links, and we can test in this way: we choose bad cases predicted by CD. For given label $y$, we select label $y^{\prime}\neq y$, and we choose the top-$3$ concepts in $\mathcal{C}_{y^{\prime}}$ as the candidate concepts to generate continuous prompts instead of submodular optimization. We call one change of concepts as “causal-attack”. And we calculate the average cosine similarities between input and generated explanation. The results are showed in Figure 4. We can infer that one possible reason of "noise issue" is the short cut link.
6.4 Longer $N_{q}$ Scenario
To test the performance when $N_{q}$ increases, we conduct experiments with $N_{q}=5$ and $N_{q}=10$ on SST-2 dataset. The results are reported in Table 5. We can draw the conclusion that CD could apply to scenarios which $N_{q}$ are larger than $1$ and $2$.
7 Conclusion
In this paper, we introduce CD, a framework to interpret continuous prompts employing generated optimized concepts to decompose the prompt embeddings. Specifically, we prove that continuous prompts could be decomposed into concepts. Then we prompt GPT-4o to generate candidates and use a submodular optimization algorithm to select the concepts. Finally we design a specific loss tune the concept embeddings and propose a local-level strategy to generate explanations. We conduct extensive experiments on three datasets across two typical language model architectures with various prompt types. Results demonstrate that the effectiveness of our framework.
Limitations
Based on the correlation experiment results, the selected concepts still suffer from noise issue, in this study we only propose a simple heuristic strategy to filter irrelevant samples. How to efficiently choose vital concepts needs further exploration.
Appendix A Proof
Lemma For any $\boldsymbol{P}\in\mathbb{R}^{d\times N_{q}}$, $\epsilon>0$, there exist $\boldsymbol{C}\in\mathbb{R}^{d\times N_{c}}$, $\boldsymbol{Q}\in\mathbb{R}^{N_{c}\times N_{q}}$ that satisfies $||\boldsymbol{CQ}-\boldsymbol{P}||_{F}^{2}\leq\epsilon$.
Proof.
First, it is straightforward that if $h(\boldsymbol{C})=\left\|\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}\right\|_{F}^{2}$ is convex, then we can obtain a $\boldsymbol{Q}^{*}$ such that $\left\|\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}\right\|_{F}^{2}\leq\epsilon$ Zinkevich (2003). Therefore we only need to prove that $h(\boldsymbol{C})$ is convex. For any $\boldsymbol{C}$,
$h(\boldsymbol{C})=\left\|\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}\right\|_{F}^{2}=tr((\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P})^{T}(\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P})).$ (12)
To prove that the function $h(\boldsymbol{C})$ is convex, we need to show that its Hessian matrix is positive semi-definite for all $\boldsymbol{C}$. Let’s denote the $ij$-th element of the Hessian matrix as $\boldsymbol{H}_{ij}$. The Hessian matrix of $h(\boldsymbol{C})$ is given by:
$\boldsymbol{H}_{ij}=\frac{\partial^{2}h}{\partial\boldsymbol{C}_{ij}\partial\boldsymbol{C}_{kl}}.$ (13)
To prove convexity, we need to show that for any vector $\boldsymbol{v}$, $\boldsymbol{v}^{T}\boldsymbol{Hv}\geq 0$.
$\boldsymbol{H}_{ij}=\frac{\partial^{2}}{\partial\boldsymbol{C}_{ij}\partial\boldsymbol{C}_{kl}}(\left\|\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}\right\|_{F}^{2}).$ (14)
Using the properties of the Frobenius norm, we can rewrite
$\begin{split}\boldsymbol{H}_{ij}&=\frac{\partial}{\partial\boldsymbol{C}_{ij}}(\frac{\partial}{\partial\boldsymbol{C}_{ij}}tr((\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P})^{T}(\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P})))\\ &=\frac{\partial}{\partial\boldsymbol{C}_{ij}}(2(\boldsymbol{Q}^{T}(\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}))_{kl})\\ &=2\frac{\partial}{\partial\boldsymbol{C}_{ij}}((\boldsymbol{Q}^{T}(\boldsymbol{C}\boldsymbol{Q}-\boldsymbol{P}))_{kl})\\ &=2((\boldsymbol{Q}^{T}\boldsymbol{Q})_{ik}\delta_{jl}),\end{split}$ (15)
$\delta_{jl}$ is the Kronecker delta. Note that $\boldsymbol{Q}^{T}\boldsymbol{Q}$ is positive semi-definite, and $\delta_{jl}\geq 0$, so $\boldsymbol{H}_{ij}$ is positive semi-definite.
Appendix B Implementation Details
In this work, all language models are implemented using Transformers. All our experiments are conducted on a single A800 GPU, and the results are reported using 5 random seeds (i.e. 1, 42, 100, 999, 1756).
For BERT-Large and GPT2-Medium checkpoints (including MedBERT, LEGAL-BERT and FinBERT), we download model weights from Huggingface Model Hub.
For Discrete-1500, due to the lack of resource code, we implement this method ourselves. Specifically, we compare the vocab embeddings with label embeddings (feeding into BERT-Large/GPT-2-Medium) and we select top-1500 tokens based on cosine similarities as our explanation vocabulary.
Appendix C Trade-off between Interpretability and Performance
We conduct an analysis on with how the number of concepts per class (i.e. $|\mathcal{C}_{y}|$) affect the accuracy performance. The results in BERT-Large ($|N_{q}|=2$) are stated in Table 7.
  Method 1 5 10 15 20
SST-2 0.691 0.653 0.745 0.637 0.539
IMDB 0.688 0.633 0.715 0.614 0.556
AGNews 0.658 0.629 0.777 0.601 0.594
Table 7: Accuracy results with variant $|\mathcal{C}_{y}|$ ($N_{q}=2$) on BERT-Large.
The results in GPT-2-Medium ($N_{q}=2$) are given in Table 8.
  Method 1 5 10 15 20
SST-2 0.652 0.667 0.730 0.619 0.547
IMDB 0.649 0.619 0.715 0.621 0.539
AGNews 0.618 0.605 0.777 0.611 0.581
Table 8: Accuracy results with variant $|\mathcal{C}_{y}|$ ($N_{q}=2$) on GPT2-Medium.
Appendix D Dataset Statistics
The statistical information for mentioned six datasets is shown in Table 9.
  Dataset Train/Dev/Test C L
SST-2 6920/872/1821 2 50
IMDB 20K/5K/25K 2 500
AGNews 25K/2K/7K 4 500
Medical Abstracts 10K/1K/2.8K 5 500
JUSTICE 2.5K/0.1K/1.8K 8 500
Finance Sentiment 3.8K/0.1K/0.9K 3 500
Table 9: Statistics of three datasets. C: number of classes, L: average text length
Appendix E Prompt List
  Prompts
1.describe what a positive movie review looks like 2.describe the aspects of a positive movie review 3.describe a positive movie review using sentences 4.describe the patterns of a positive movie review 5.describe the concepts of a positive movie review 6.describe what a negative movie review looks like 7.describe the aspects of a negative movie review 8.describe a negative movie review using sentences 9.describe the patterns of a negative movie review 10.describe the concepts of a negative movie review
Table 10: Prompts to generate candidate concepts on SST-2 dataset.
  Prompts
1.describe what a positive review looks like 2.describe the aspects of a positive review 3.describe a positive review using sentences 4.describe the patterns of a positive review 5.describe the concepts of a positive review 6.describe what a negative review looks like 7.describe the aspects of a negative review 8.describe a negative review using sentences 9.describe the patterns of a negative review 10.describe the concepts of a negative review
Table 11: Prompts to generate candidate concepts on IMDB dataset.
  Prompts
1.describe what a world news looks like 2.describe the aspects of a world news 3.describe what a sports news looks like 4.describe the aspects of a sports news 5.describe what a business news looks like 6.describe the aspects of a business news 7.describe what a sci/tech news looks like 8.describe the aspects of a sci/tech news
Table 12: Prompts to generate candidate concepts on AGNews dataset.
  Prompts
1.use some phrases to describe neoplasms 2.use some sentences to describe neoplasms 3.use some phrases to describe digestive system diseases 4.use some sentences to describe digestive system diseases 5.use some phrases to describe nervous system diseases 6.use some sentences to describe nervous system diseases 7.use some phrases to describe cardiovascular disease 8.use some sentences to describe cardiovascular disease 9.use some phrases to describe general pathological conditions
Table 13: Prompts to generate candidate concepts on Medical Abstracts dataset.
  Prompts
1.use some phrases to describe majority opinion in legal 2.use some phrases to describe per curiam in legal terms 3.use some phrases to describe plurality opinion in legal terms 4.use some phrases to describe equally divided in legal terms 5.use some phrases to describe dismissal-moot in legal terms 6. use some phrases to describe dismissal-improvidently-granted in legal terms 7. use some phrases to describe dismissal-other in legal terms 8.use some phrases to describe opinion of the court in legal terms
Table 14: Prompts to generate candidate concepts on Justice dataset.
  Prompts
1.describe the aspects of a finance positive review 2.what a finance positive review may include 3.describe the aspects of a finance neutral review 4.what a finance neutral review may include 5.describe the aspects of a finance negative review 6.what a finance negative review may include
Table 15: Prompts to generate candidate concepts on Finance Sentiment dataset.
  Prompts
1.[$x$]The feeling of the review is [$p$][$m$]. 2.[$x$]The feeling of the review is [$p$][$p$][$m$]. 3.[$x$]What emotion does the review evoke ? [$p$][$m$] 4.[$x$]What emotion does the review evoke ? [$p$][$p$][$m$] 5.[$x$]describe the sentiment of the text. [$p$][$m$] 6.[$x$]describe the sentiment of the text. [$p$][$p$][$m$]
Table 16: Prompts used in Table2 on SST-2, IMDB dataset. $m$ is the mask token. $p$ is the prefix token.
Appendix F More Prompt Results
BERT-Large GPT-2-Medium
$N_{q}=1$ $N_{q}=2$ $N_{q}=1$ $N_{q}=2$
P-tuning CD P-tuning CD P-tuning CD P-tuning CD
  What emotion does the review evoke 4-shot 0.527 0.518 0.568 0.560 0.534 0.531 0.547 0.534
8-shot 0.562 0.547 0.581 0.676 0.559 0.577 0.596 0.614
16-shot 0.781 0.773 0.744 0.744 0.807 0.813 0.747 0.725
32-shot 0.830 0.840 0.794 0.812 0.843 0.843 0.825 0.819
Full 0.898 0.907 0.893 0.892 0.909 0.911 0.864 0.902
Avg 0.719 0.717 0.751 0.736 0.745 0.735 0.715 0.718
${\sigma^{2}}$${\pm 0.003}$${\pm 0.001}$${\pm 0.001}$$\pm 0.003$$\pm 0.003$$\pm 0.005$$\pm 0.0003$$\pm 0.002$
  describe the sentiment of the text 4-shot 0.545 0.592 0.575 0.592 0.549 0.545 0.548 0.542
8-shot 0.557 0.615 0.576 0.615 0.582 0.593 0.571 0.636
16-shot 0.718 0.797 0.730 0.797 0.814 0.803 0.771 0.787
32-shot 0.836 0.850 0.821 0.850 0.817 0.832 0.818 0.843
Full 0.896 0.907 0.895 0.907 0.908 0.909 0.887 0.891
Avg 0.710 0.752 0.719 0.752 0.734 0.736 0.719 0.739
$\sigma^{2}$$\pm 0.0007$$\pm 0.005$$\pm 0.003$$\pm 0.005$$\pm 0.003$$\pm 0.0002$$\pm 0.0004$$\pm 0.00006$
Table 17: Accuracy results of CD with various prompt types on the SST-2 dataset. $N_{q}$ denotes the number of continuous prompt tokens. $\sigma^{2}$ denotes the average variance across five shot settings.
BERT-Large GPT-2-Medium
$N_{q}=1$ $N_{q}=2$ $N_{q}=1$ $N_{q}=2$
P-tuning CD P-tuning CD P-tuning CD P-tuning CD
  What emotion does the review evoke 4-shot 0.551 0.559 0.543 0.543 0.534 0.531 0.519 0.525
8-shot 0.545 0.546 0.542 0.532 0.559 0.577 0.502 0.500
16-shot 0.579 0.549 0.693 0.663 0.807 0.813 0.501 0.502
32-shot 0.766 0.823 0.818 0.828 0.843 0.843 0.700 0.640
Full 0.908 0.910 0.914 0.914 0.918 0.921 0.902 0.914
Avg 0.669 0.677 0.702 0.696 0.732 0.737 0.624 0.616
$\sigma^{2}$$\pm 0.0009$$\pm 0.0005$$\pm 0.0012$$\pm 0.0014$$\pm 0.0003$$\pm 0.0004$$\pm 0.0009$$\pm 0.0012$
  describe the sentiment of the text 4-shot 0.549 0.542 0.541 0.592 0.551 0.549 0.522 0.508
8-shot 0.548 0.528 0.525 0.615 0.517 0.510 0.513 0.501
16-shot 0.660 0.620 0.660 0.797 0.503 0.501 0.503 0.501
32-shot 0.709 0.698 0.799 0.850 0.714 0.632 0.594 0.621
Full 0.915 0.911 0.919 0.919 0.917 0.927 0.900 0.908
Avg 0.676 0.659 0.688 0.754 0.640 0.623 0.606 0.607
$\sigma^{2}$$\pm 0.0021$$\pm 0.0014$$\pm 0.0009$$\pm 0.0016$$\pm 0.0004$$\pm 0.0002$$\pm 0.0007$$\pm 0.0003$
Table 18: Accuracy results of CD with various prompt types on the IMDB dataset. $N_{q}$ denotes the number of continuous prompt tokens. $\sigma^{2}$ denotes the average variance across five shot settings.
BERT-Large GPT-2-Medium
$N_{q}=1$ $N_{q}=2$ $N_{q}=1$ $N_{q}=2$
P-tuning CD P-tuning CD P-tuning CD P-tuning CD
  The primary focus of this news is 4-shot 0.733 0.733 0.728 0.736 0.746 0.757 0.759 0.758
8-shot 0.771 0.732 0.779 0.766 0.785 0.776 0.791 0.795
16-shot 0.800 0.826 0.793 0.791 0.788 0.789 0.793 0.798
32-shot 0.841 0.844 0.844 0.842 0.825 0.824 0.834 0.828
Full 0.885 0.886 0.874 0.877 0.863 0.867 0.859 0.858
Avg 0.806 0.804 0.822 0.802 0.801 0.802 0.807 0.807
$\sigma^{2}$$\pm 0.0004$$\pm 0.0004$$\pm 0.0001$$\pm 0.0003$$\pm 0.0001$$\pm 0.0001$$\pm 0.0004$$\pm 0.0006$
  The core theme of this news is 4-shot 0.697 0.777 0.652 0.728 0.778 0.771 0.765 0.778
8-shot 0.788 0.752 0.796 0.768 0.795 0.790 0.788 0.793
16-shot 0.806 0.795 0.810 0.819 0.791 0.801 0.789 0.789
32-shot 0.831 0.831 0.829 0.839 0.834 0.840 0.825 0.831
Full 0.886 0.888 0.876 0.875 0.867 0.865 0.853 0.863
Avg 0.801 0.808 0.792 0.805 0.813 0.813 0.804 0.810
$\sigma^{2}$$\pm 0.0002$$\pm 0.0003$$\pm 0.0005$$\pm 0.0004$$\pm 0.0002$$\pm 0.0005$$\pm 0.0001$$\pm 0.0006$
Table 19: Accuracy results of CD with various prompt types on the AGNews dataset. $N_{q}$ denotes the number of continuous prompt tokens. $\sigma^{2}$ denotes the average variance across five shot settings.